# Umberto Collodel Central Bank of Malta

### v1

@@ Abstract
This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using a Large Language Model (LLM). We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles. These agents forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with the variation across forecasts serving as a measure of market uncertainty or disagreement. We evaluate three prompting strategies, naive, few-shot (enriched with historical data), and an advanced iterative “LLM-as-a-Judge” framework, to assess the effect of prompt design on predictive performance. Even the naive approach generates a strong correlation (~0.5) between synthetic disagreement and actual market outcomes, particularly for longer-term maturities. The LLM-as-a-Judge framework further improves accuracy at the first iteration. These results demonstrate that LLM-driven simulations can capture interpretive uncertainty beyond traditional measures, providing central banks with a practical tool to anticipate market reactions, refine communication strategies, and enhance financial stability.
@@ 1 Introduction
Recent advancements in large language models (LLMs) have opened up a plethora of possibilities for economic and financial research. Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior. Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023). Notwithstanding their growing capabilities, the application of LLMs in finance has so far been relatively limited. [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023). Yet, the financial domain may be one of the most promising areas for their use. Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings. [^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009). In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011). However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy. It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents. Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019). In previous research, we moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025).333(Bauer et al.,, 2021) find similar results for the US. Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
@@
To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies. The first is a naive or ”zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style. The second is a ”few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them. Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023). In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score. The Judge then rewrites the prompt to improve alignment with market reactions. This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations. The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets. Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework, significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter. This breakthrough offers policymakers a powerful new tool to preemptively evaluate and refine central bank communications, identifying language that may trigger market disagreement and volatility well before public release. By moving beyond traditional sentiment analysis, our approach provides a nuanced understanding of how monetary policy signals are perceived and transmitted through financial markets. Integrating insights from macroeconomic communication, behavioral finance, and artificial intelligence, this framework presents a transformative lens for both scholars and decision-makers to better anticipate and manage the financial market impact of monetary policy announcements.
@@ 2 Literature Review
The paper contributes to three intersecting literatures: monetary policy communication, behavioral expectations, and the emerging use of large language models (LLMs) in economic analysis.
@@
A second application is the construction of LLM-powered simulated economic agents, “homo silicus” (Horton,, 2023)—who are endowed with preferences, information, and decision rules, and can interact in artificial market environments. Frameworks like EconAgent (Zhang et al.,, 2023) and ASFM (Li et al.,, 2024) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
@@
Our approach delivers a methodological advance with clear policy relevance. We provide the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions. This design closes a long-standing gap between the empirical identification of communication shocks and the theoretical modeling of expectation formation. By combining sentiment extraction, agent-based modeling, and LLM-driven language interpretation in a unified environment, we move beyond reduced-form event studies and open the “black box” of how policy language is processed by market participants. The framework is adaptable across institutional settings, enabling comparative analyses of communication strategies by central banks worldwide. This makes it directly relevant to current debates on transparency, forward guidance, and the role of central bank credibility in shaping financial conditions in both advanced and emerging market economies.
3 Methodology
@@
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on European Central Bank (ECB) press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between 1998-06-09 and 2025-04-17.
@@
Figure 1 illustrates market-based disagreement following ECB Governing Council meetings across all three maturities. The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent tightening cycle initiated in 2022. Volatility exhibits a clear maturity structure, with consistently higher levels at longer tenors reflecting greater uncertainty about the medium- and long-term policy trajectory compared to near-term rate expectations.
@@ 3.1 Agents
We construct a population of 30 synthetic agents, each instantiated through Gemini 2.5-Flash model, one of Google’s most recent large language model. [^6] We use 30 agents to balance computational and cost efficiency with sufficient heterogeneity to capture cross-sectional variation. The agents are designed to interpret the natural language of ECB press conferences and produce quantitative forecasts of Euro IRS rates at the three maturities. The simulation proceeds as follows. Each agent is endowed with:
@@
    Risk aversion parameter – fixed at the start of the simulation, drawn from a pre-defined distribution to induce heterogeneity in risk preferences.
@@
    Cognitive bias profile – encompassing tendencies such as optimism, pessimism, anchoring, or overreaction, assigned to influence interpretation.
@@
    Interpretive style – a qualitative descriptor shaping the reading of ECB communication (e.g., focusing on inflation outlook, growth projections, or geopolitical risks).
These attributes are embedded in the prompt given to the LLM for each agent, ensuring that the same press conference produces different interpretations across agents. [^7] We set the temperature parameter equal to 1. Temperature is a parameter that controls the randomness of a language model’s output. Lower values (e.g., 0.2) make the model more deterministic and focused, while higher values (e.g., 1.0 or above) increase creativity and variability by making the model more likely to sample less probable words.
@@ 3.2 Prompting Strategies
The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes.
@@
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions. [^8] Big LLMs are notoriously difficult to fine-tune because of computational and cost-related issues. Moreover, we do not have enough observations for this use case. In these instances, a ”few-shot” technique has been shown to improve over a ”zero-shot” benchmark (Brown et al.,, 2020).
@@
Introduces a second LLM, the “Judge,” which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment. The improved prompt is fed back into the “Analyst” LLM, forming an adaptive feedback loop. The process continues over multiple iterations before all the prompts provided by the analyst are tested out-of-sample. [^9] The split train-test is 75/25 in chronological order. [^10] Following the literature, we use for the ”Judge” a more potent model compared to the ”Analyst”, namely Gemini 2.5-Pro [^11] See Algorithm 1 for the detailed implementation steps. [^12] See Figure 11 for the judge prompt.
@@ 3.3 Evaluation
We evaluate the ability of the LLM to reproduce markets’ heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data. Our main metric is the Spearman correlation coefficient, computed separately for short, medium, and long maturities. We also compare in-sample and out-of-sample performance for each prompting strategy in the LLM-as-Judge approach to gauge the extent of overfitting. [^15] The LLM-as-a-Judge approach is expected to produce higher out-of-sample correlations if the adaptive prompt optimization captures persistent features of agent interpretation rather than idiosyncratic event noise. Finally, we analyze whether certain maturities are more sensitive to synthetic disagreement, providing insights into which parts of the yield curve are most affected by interpretive heterogeneity in response to central bank communication. Figure 3 shows a stylized illustration of the whole process. [^16] For each press conference, we send a single API request. Although batching multiple conferences would significantly reduce computation time—since the request itself is the slowest part of the process—we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.

### v2

@@ Abstract
This paper develops a novel method to simulate financial market reactions to European Central Bank (ECB) press conferences using Large Language Models (LLMs). We create a behavioral, agent-based simulation of 30 synthetic traders, each with distinct risk preferences, cognitive biases, and interpretive styles. These agents interpret ECB communication and forecast Euro interest rate swap levels at 3-month, 2-year, and 10-year maturities, with cross-sectional variation serving as a measure of market uncertainty. Using a comprehensive dataset of 283 ECB press conferences (1998-2025), we evaluate three prompting strategies: naive zero-shot, few-shot (enriched with historical context), and an advanced iterative ”LLM-as-a-Judge” framework. We find that even the naive approach achieves substantial correlations (roughly 0.5) between synthetic disagreement and actual post-announcement OIS volatility, particularly for medium- and long-term maturities. The LLM-as-a-Judge framework further enhances performance upon initial application, reaching correlations of almost 0.6 and identifying key drivers of volatility. Our results demonstrate that LLM-based agents can effectively capture the heterogeneous interpretive processes through which monetary policy signals propagate to financial markets, moving beyond traditional reduced-form event studies. This provides central banks with a practical tool to ex-ante evaluate communication strategies and anticipate market reactions, while offering researchers a micro-founded framework for understanding expectation formation in response to central bank communication.
@@ 1 Introduction
Recent advancements in Large Language Models (LLMs) have opened up a plethora of possibilities for economic and financial research. Due to the way these models are trained, on vast and diverse textual data, and their ability to generate contextually rich and realistic responses, LLMs can now be viewed as implicit computational representations of human behavior. Much like the economic construct of homo economicus, these models can be endowed with information, placed in hypothetical scenarios, and observed for their responses (Horton,, 2023; Kazinnik and Sinclair,, 2025). Despite their growing capabilities, the application of LLMs in finance has been relatively limited so far. [^1] A notable exception is, for example, Lopez-Lira and Tang, (2023). Yet, the financial domain may be one of the most promising areas for their use. Unlike traditional economic models that assume rational agents and efficient markets, real-world financial markets are often driven by sentiment, narrative, and behavioral biases, factors that are difficult to capture with purely mathematical models, especially in high-frequency settings. [^2] Agent-based models (ABMs) have emerged as a flexible framework to incorporate such heterogeneity (Farmer and Foley,, 2009). In the context of monetary policy, ABMs have been used to simulate expectation formation, market volatility, and policy transmission (Delli Gatti et al.,, 2011). However, most ABMs rely on stylized behavioral rules imposed by the researcher (Horton,, 2023).
A key area where LLMs can offer significant value is in modeling reactions to central bank communication, a cornerstone of modern monetary policy. It is well-established that the effectiveness of monetary policy hinges not just on the decisions themselves, but on the language used and its interpretation by economic agents. Press conferences, speeches, and official statements play a pivotal role in shaping market expectations and, consequently, asset prices (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019). Previous research, has moved beyond the general direction of the market response to highlight the crucial role of response heterogeneity; in the Euro area, high disagreement following a statement transmits to financial markets through a separate channel compared to the standard monetary policy surprise channel, and can dampen the intended impact of policy tightening by lowering the term premium of the yield curve (Collodel and Kunzmann,, 2025). [^3] Bauer et al., (2021) find similar results for the US. Capturing this nuanced, belief-driven heterogeneity is a challenge for traditional models but is a natural fit for the capabilities of LLMs.
@@
To further explore how the design of prompting influences this correlation, we implement three distinct prompting strategies. The first is a naive or “zero-shot” prompt (Radford et al.,, 2019), which serves as the baseline: each agent is assigned a simple, pre-defined level of risk aversion, a set of cognitive biases, and a unique interpretive style. The second is a “few-shot” prompt (Brown et al.,, 2020), which enriches the agent’s context by including the past three ECB press conferences along with the corresponding changes in market-based disagreement resulting from them. Finally, we introduce an LLM-as-a-Judge prompt, our most advanced strategy (Zheng et al.,, 2023). In this iterative setup, a separate LLM, the Judge, acts as a critical evaluator, systematically reviewing the analyst prompt and its resulting in-sample correlation score. The Judge then rewrites the prompt to improve alignment with market reactions. This revised prompt is fed back into the Analyst LLM, forming a feedback loop that continues over different iterations. The different prompt versions are subsequently evaluated out-of-sample to assess their robustness and decrease concerns relative to model over-fitting.
We find that LLM-based agents can effectively replicate the complex and varied reactions of investors to ECB press conferences, achieving a remarkable correlation of around 0.5 between the synthetic disagreement measure and actual market outcomes for medium- and long-term maturities, a substantial result given the inherent complexity and noise of financial markets. Contrary to expectations, a more detailed few-shot prompt does not improve this correlation; however, the LLM-as-a-Judge framework significantly enhances predictive accuracy upon its initial application, with performance degrading thereafter. These results suggest that LLM-based agent simulations may offer policymakers a useful tool for evaluating central bank communications and identifying language that could generate market disagreement prior to public release. By moving beyond traditional sentiment analysis, our approach contributes to understanding how monetary policy signals are perceived and transmitted through financial markets. The framework integrates insights from macroeconomic communication, behavioral finance, and artificial intelligence to help scholars and policymakers better understand the financial market impact of monetary policy announcements.
The remainder of the paper proceeds as follows. Section 2 reviews the relevant literature on monetary policy communication, behavioral expectations, and LLM applications in economics. Section 3 outlines our methodology, including data collection of ECB press conferences and market reactions, construction of synthetic agent personas, and our three prompting strategies. Section 4 presents our empirical results, analyzing agent forecast dispersion patterns, correlation with market volatility measures, and comparative performance across prompting approaches. Section 5 probes the robustness of our results to model and prompt variation. Lastly, Section 6 concludes with policy implications for central bank communication strategies and directions for future research.
@@ 2 Literature Review
The paper contributes to three intersecting strands of economic literature: monetary policy communication, behavioral expectations, and the emerging use of LLMs in economic analysis.
@@
A second application is the construction of LLM-powered simulated economic agents, homo silicus (Horton,, 2023; Kazinnik and Sinclair,, 2025), who are endowed with preferences, information, and decision rules, and can interact in artificial market environments. Frameworks like EconAgent (Zhang et al.,, 2023) produce heterogeneous, adaptive behavior that is richer and more realistic than the stylized rules of conventional ABMs, while also capturing the interpretation of complex qualitative signals.
@@
Our approach delivers a methodological advance with clear policy relevance. Specifically, it provides the first micro-founded framework in which heterogeneous, boundedly rational agents interpret central bank communication in real time and translate qualitative policy signals into market-relevant actions. This design closes a long-standing gap between the empirical identification of communication shocks and the theoretical modeling of expectation formation. By combining sentiment extraction, agent-based modeling, and LLM-driven language interpretation in a unified environment, we move beyond reduced-form event studies and open the “black box” of how policy language is processed by market participants. The framework is adaptable across institutional settings, enabling comparative analyses of communication strategies by central banks worldwide. This makes it directly relevant to current debates on transparency, forward guidance, and the role of central bank credibility in shaping financial conditions in both advanced and emerging market economies.
3 Methodology and Data
@@
Our methodology proceeds in three main steps: (i) construction of a structured database of ECB communication events and corresponding market reactions; (ii) development of LLM-based synthetic agents embedded in an agent-based simulation framework; and (iii) evaluation of the correspondence between synthetic measures of belief heterogeneity and actual market outcomes. We focus on ECB press conferences, which constitute the primary vehicle for conveying monetary policy decisions in the Euro area. Following the literature on high-frequency identification of monetary policy shocks (Gurkaynak et al.,, 2005; Altavilla et al.,, 2019), we define event dates as the days of regularly scheduled ECB Governing Council monetary policy meetings, limiting the sample to those that include both a written policy statement and a subsequent press conference by the ECB President. For each event, we collect the full transcript of the press conference as published on the ECB website as well as the associated OIS post-conference volatility for the short (3-months), medium (2-years), and long (10-years) term. [^5] Following the monetary policy communication literature, we use OIS rates as they reflect pure interest rate expectations without the credit risk component present in government bond yields (Gurkaynak et al.,, 2005). Our final sample comprises 283 ECB conferences between June 1998 and April 2025.
@@
Figure 1 illustrates market-based disagreement following ECB Governing Council meetings across all three maturities. The data reveal pronounced spikes during periods of heightened policy uncertainty, notably the 2008-2009 financial crisis, the European sovereign debt crisis, and the recent monetary policy tightening cycle initiated in 2022. Volatility exhibits a clear maturity structure, with consistently higher levels at longer tenors reflecting greater uncertainty about the medium- and long-term policy trajectory compared to near-term rate expectations.
@@ 3.1 Agents
We construct a population of 30 synthetic agents, each instantiated through Google’s Gemini 2.5-Flash model. [^6] We select 30 agents to balance computational efficiency with sufficient heterogeneity to capture meaningful cross-sectional variation while maintaining tractable inference costs. Sensitivity analysis with 20 and 50 agents yields qualitatively similar results. The agents are designed to interpret ECB press conference transcripts and generate quantitative forecasts of Euro IOS rates across three key maturities: 3-month, 2-year, and 10-year tenors. Each agent is endowed with distinct characteristics to ensure interpretive heterogeneity:
@@
    Risk Aversion Profile: Categorical assignment (High/Medium/Low) determining sensitivity to uncertainty and preference for yield curve stability.
@@
    Cognitive Bias Portfolio: Each agent receives 1-2 behavioral biases drawn from established behavioral finance literature, including confirmation bias, overconfidence, anchoring effects, herding tendencies, loss aversion, and recency bias. This design reflects empirical evidence that real market participants exhibit systematic deviations from rational expectations.
@@
    Interpretive Framework: Agents adopt distinct analytical lenses (Fundamentalist, Technical, Sentiment-driven, Quantitative, Narrative-focused, or Policy-skeptic) that shape their processing of ECB communication content.
The LLM dynamically assigns these characteristics to create 30 heterogeneous agents for each press conference, ensuring that identical content generates diverse interpretations: this approach allows the model to adapt agent composition naturally over time while preserving the behavioral diversity necessary to capture cross-sectional variation in market reactions. [^7] We set the temperature parameter equal to 1. Temperature is a parameter that controls the randomness of a language model’s output. Lower values (e.g., 0.2) make the model more deterministic and focused, while higher values (e.g., 1.0 or above) increase creativity and variability by making the model more likely to sample less probable words.
@@ 3.2 Prompting Strategies
The baseline approach, where the agent receives only the ECB press conference transcript along with its fixed attributes. This approach tests whether LLMs possess intrinsic capabilities to decode monetary policy communication without explicit calibration.
@@
Augments the baseline by including the transcripts of the three most recent ECB press conferences and the associated historical changes in market-based disagreement, providing a short memory of past communication–market interactions.This design tests whether contextual examples improve forecasting accuracy, following established few-shot learning principles. [^8] Few-shot prompting has demonstrated superior performance over zero-shot approaches across multiple economic forecasting applications, particularly when fine-tuning is computationally infeasible.
@@
Introduces a second LLM, the “Judge”, which reviews each agent’s initial prompt and in-sample performance (correlation with market disagreement) and iteratively rewrites the prompt to improve alignment. [^9] Following the literature, we use for the “Judge” a more potent model compared to the “Analyst”, namely Gemini 2.5-Pro. The improved prompt is fed back into the “Analyst” LLM, forming an adaptive feedback loop. The process continues over multiple iterations before all the prompts provided by the analyst are tested out-of-sample. This approach enables dynamic prompt optimization while maintaining rigorous out-of-sample validation to address overfitting concerns. [^10] The split train-test is 75/25 in chronological order. [^11] See Algorithm 1 for the detailed algorithm and Figure 17 for the “Judge” prompt.
@@ 3.3 Evaluation
We evaluate the ability of the LLM to reproduce market heterogeneity by correlating the synthetic disagreement measures with the empirical disagreement observed in market data. Our main metric is the Spearman correlation coefficient, computed separately for each maturity. We employ Spearman correlation as our primary evaluation metric for several methodological reasons. First, Spearman correlation captures monotonic relationships without imposing linearity assumptions, which is particularly appropriate given the potentially complex and non-linear mapping between textual ambiguity in central bank communication and subsequent market volatility. Second, it demonstrates superior robustness to outliers and extreme events that frequently characterize financial market data, especially during crisis periods. Nevertheless, to ensure our findings are not dependent on this methodological choice, we conduct robustness checks using alternative correlation metrics. [^14] See Section Appendix B: LLM-as-Judge Implementation Details and Results. We also compare in-sample and out-of-sample performance for each prompting strategy in the LLM-as-Judge approach to gauge the extent of overfitting. [^15] The LLM-as-a-Judge approach is expected to produce higher out-of-sample correlations if the adaptive prompt optimization captures persistent features of agent interpretation rather than idiosyncratic event noise. Finally, we analyze whether certain maturities are more sensitive to synthetic disagreement, providing insights into which parts of the yield curve are most affected by interpretive heterogeneity in response to central bank communication. Figure 3 shows a stylized illustration of the whole process. [^16] For each press conference, we send a single API request. Although batching multiple conferences would significantly reduce computation time, since the request itself is the slowest part of the process, we adopt a conservative approach to avoid including future information as context, which could inadvertently influence the model’s output.
@@ 4 Results
This section evaluates the performance of our three prompting strategies in replicating market disagreement patterns following ECB press conferences. We examine the zero-shot baseline (Section 4.1), the few-shot approach with historical context (Section 4.2), and the iterative LLM-as-a-Judge framework (Section 4.3), focusing on correlations between synthetic disagreement and actual market volatility across different yield curve tenors. We conclude with a comparison against simple textual benchmarks (Section LABEL:sec:complexity) to demonstrate the value of modeling heterogeneous agent interpretations compared to less sophisticated approaches.